Oz Jeans ha factories in Mydney and Selbourne. At the Mydney factory fixed costs are $28000 per month and the cost of producing each pair of jeans is $30. At the Selbourne factory, fixed costs are $35200 per month and each pair of jeans costs $24 to produce. During the next month Oz Jeans must manufacture 6000 pairs of jeans. Calculate the production order for each factory, if the total manufacturing costs for each factory are to be the same.
Answer by stanbon(75887)

At the Mydney factory fixed costs are $28,000 per month and the cost of producing each pair of jeans is $30.
At the Selbourne factory, fixed costs are $35,200 per month and each pair of jeans costs $24 to produce.
During the next month Oz Jeans must manufacture 6000 pairs of jeans.
Calculate the production order for each factory, if the total manufacturing costs for each factory are to be the same.
-----------------------
Let "x" be number of units to be produced:
-----------------------
Mydney costs: M(x)=28000+30x
Selbourne costs: S(x)=35200+24(6000-x)
-----------------------
EQUATION:
28000+30x = 35200+24(6000-x)
28000+30x = 35200 + 144000-24x
54x = 151200
x = 2800 (# of jeans to be produced by Mydney)
6000-x = 3200 (# of jeans to be produced by Selbourne)
================
Cheers,
Stan H.